# The Cultural Gene of Large Language Models: A Study on the Impact of Cross-Corpus Training on Model Values and Biases

### v3

@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) represents a paradigm shift in artificial intelligence, with systems demonstrating remarkable capabilities in understanding, generating, and interacting with[Besta2023GraphofThoughts, 10090421] human language . This technological leap has catalyzed the integration of AI into the very fabric of our digital and physical lives. AI systems now optimize complex, heterogeneous edge networks through novel strategies like experience-driven model migration [10508334], and they enhance the efficiency of federated learning via sophisticated neural architecture searches . The reach of AI extends even into our physical environment, with novel sensing technologies leveraging commodity WiFi for applications ranging from gesture recognition [10090421] and anti-interference activity detection [9613773, laskin2022context] to unobtrusive emotion recognition [10149418, hoffmann2022training] and even respiratory healthcare monitoring [10804189, penedo2023decoupling]. This pervasive integration underscores the profound impact of modern machine learning, built upon powerful techniques like contrastive learning [liu2024deja, gu2023mamba] and efficient model compression .
@@
This paper posits that LLMs trained on culturally-specific corpora inherit what we term a ”cultural gene”—a deeply embedded, systematic tendency to reason and respond in ways that align with the values and norms of that culture. This concept moves beyond surface-level biases in language generation [sheng2019woman] or the presence of stereotypes [nadeem-etal-2021-stereoset], and instead seeks to uncover the foundational worldview imprinted upon the model. The foundational work of Hofstede on cultural dimensions provides a powerful framework for conceptualizing these differences, such as the spectrum between individualism and collectivism or variations in power distance [hofstede1980cultures]. Recent work has begun to probe these dimensions, asking whether models possess cultural intelligence [bleu2023language] or how they represent knowledge from different cultures [jiang2023large], but a systematic methodology for quantifying these ingrained ”genes” is still nascent. The challenge is immense, as these models are increasingly tasked with sensitive roles in domains like facial expression recognition, where cultural display rules are paramount and label noise can be a significant confounder [chiang2023vicuna, 10916982].
The stakes of failing to address this challenge are substantial. Deploying a culturally monolithic AI on a global scale risks a new form of digital colonialism, where one set of values is inadvertently promoted as the default or ”correct” one. This can lead to misunderstandings, perpetuate harmful stereotypes [mhamdi-etal-2023-cross], and create systems that are ineffective or even offensive in different cultural contexts. The problem is exacerbated in multimodal systems, where visual and linguistic cues are intertwined, such as in visual dialog or video grounding , and in security applications where understanding human behavior is key, from recognizing micro-actions [ouyang2022training] to detecting malicious intent in side-channel attacks [bai2022training] or against authentication protocols [10339891, zhang2022opt].
@@ 2.1 Bias and Fairness in Language Models
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [10916982]. Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 2.2 Cross-Cultural and Multilingual NLP
The globalization of AI has spurred a growing interest in cross-cultural and multilingual NLP. Researchers have begun to move beyond English-centric benchmarks [raji2021ai], developing resources for evaluating commonsense reasoning in multiple languages [lin-etal-2021-common]. This has revealed significant disparities, highlighting the poor performance of mainstream models on tasks centered around non-Western cultures, such as those in the Africentric NLP space [ogueji2021state]. The very notion of ”cultural intelligence”[cui2025learningheterogeneitygeneralizingdynamic] in LLMs is now being formally investigated [bleu2023language], with studies directly probing models for cross-cultural differences in values [arora-etal-2023-probing].
@@ 2.3 Probing and Interpreting Model Knowledge
Researchers are now pushing beyond factual knowledge to probe more abstract concepts. Studies have analyzed how models represent social relations and roles [artetxe-etal-2023-representation], and groundbreaking work is using the geometry of a model’s internal activation space to uncover its true ”beliefs,” independent of its textual output [marks2023geometry]. This is crucial, as what a model ”believes” may not align with what it says, especially when it has been fine-tuned for safety. Analyzing the internal mechanisms, such as the structure of attention heads [vig2019analyzing] or using causal analysis to understand syntactic agreement [vig-etal-2020-causal], provides a deeper view into how these representations are formed. The ultimate goal is to understand not just *that* a model holds certain cultural values, but *why* and *how* it applies them, a question that is now being explored in the context of social norms [zhu2024can]. This deep level of analysis is essential for building trustworthy AI, whether it is used for video grounding or for complex, decentralized tasks like federated learning in edge networks [10508334].
@@ 2.4 Value Alignment and AI Ethics
This has led to calls for a more critical, decolonizing perspective on digital technologies [mhlambi2023decolonising] and the development of new evaluation frameworks that go beyond standard accuracy metrics. Researchers are now assessing LLMs against international human rights norms [shelby2023norms], testing them for social intelligence [sap-etal-2019-social], and even evaluating their ”personality” using standardized psychological tests [miotto2024do]. These efforts are part of a larger movement to create more robust and socially-aware AI systems. This is critical as AI is deployed in high-stakes, human-centric applications, from developing more efficient model compression techniques and robust neural architecture search methods [10251628], to securing our digital lives against sophisticated attacks on authentication systems [10339891] or privacy via side-channels [hoffmann2022training]. Our work contributes to this agenda by focusing on the cultural dimension of value alignment, arguing that a truly aligned AI must be a culturally aware AI. While benchmarks for Polish politics [kocon2023arrabest] or heterogeneous information retrieval [thakur2021beir] are steps in the right direction, a systematic methodology for comparing foundational cultural values is a critical, and currently missing, piece of the puzzle.
@@ 3.1 Bias and Fairness in Language Models
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones [10916982]. Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 5.1.2 Cultural Alignment Index (CAI)
The Cultural Alignment Index (CAI), as defined in Equation LABEL:eq:cai, measures how closely each model’s performance aligns with the established Hofstede scores for the USA (a high-IDV, low-PDI country) and China (a low-IDV, high-PDI country). The results are unambiguous. For the IDV dimension, $M_{W}$’s alignment with the USA was 0.91, while its alignment with China was only 0.45. The pattern was reversed for $M_{E}$, which showed an alignment of 0.85 with China and 0.48 with the USA. The same cross-alignment was observed for the PDI dimension.
@@ 5.4.2 Implications for Global AI Deployment and Ethics
Therefore, a one-size-fits-all approach to value alignment is untenable [hendrycks2021aligning]. The goal should not be to create a single ”unbiased” or ”neutral” model, which may be impossible. Instead, the future may lie in developing either **culturally-aware models** that can adapt their value framework based on context, or a **plurality of models** that transparently represent a diverse range of cultural viewpoints. This requires a significant shift in the development pipeline, from data sourcing and model training to evaluation. The need for robust systems that can handle heterogeneity is a common thread in modern computing, from federated learning with non-IID data [10508334] to wireless systems that must operate in noisy, unpredictable environments [9613773, 10339891].
@@ 5.4.3 Limitations and Future Work
Our study, while providing strong evidence, has several limitations that open avenues for future research. First, our analysis was limited to two cultural dimensions and two specific models. Future work should expand this framework to include more dimensions (e.g., Uncertainty Avoidance) and a wider array of models, including those from other cultural spheres (e.g., India, the Middle East). Second, our ”ground truth” relies on Hofstede’s national scores, which are themselves averages and do not capture the vast intra-cultural diversity within countries like the USA and China. Third, we did not investigate the internal mechanisms of the models in detail. Future work could use techniques from mechanistic interpretability [marks2023geometry, vig-etal-2020-causal] to pinpoint exactly where and how these cultural representations are stored in the network. Finally, our study is diagnostic. The next critical step is prescriptive: developing methods to effectively steer or mitigate these ingrained cultural biases, moving beyond simple debiasing techniques [gonen-goldberg-2019-lipstick] to more fundamental interventions in the training and fine-tuning process. This is a grand challenge, akin to building systems that are robust not just to data noise [10916982] but to the very fabric of cultural bias woven into human language.

### v4

@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) represents a paradigm shift in artificial intelligence, with systems demonstrating remarkable capabilities in understanding, generating, and interacting with[Besta2023GraphofThoughts] human language . This technological leap has catalyzed the integration of AI into the very fabric of our digital and physical lives. AI systems now optimize complex, heterogeneous edge networks through novel strategies like experience-driven model migration , and they enhance the efficiency of federated learning via sophisticated neural architecture searches . The reach of AI extends even into our physical environment, with novel sensing technologies leveraging commodity WiFi for applications ranging from gesture recognition and anti-interference activity detection [laskin2022context] to unobtrusive emotion recognition [hoffmann2022training] and even respiratory healthcare monitoring [penedo2023decoupling]. This pervasive integration underscores the profound impact of modern machine learning, built upon powerful techniques like contrastive learning [liu2024deja, gu2023mamba] and efficient model compression .
@@
This paper posits that LLMs trained on culturally-specific corpora inherit what we term a ”cultural gene”—a deeply embedded, systematic tendency to reason and respond in ways that align with the values and norms of that culture. This concept moves beyond surface-level biases in language generation [sheng2019woman] or the presence of stereotypes [nadeem-etal-2021-stereoset], and instead seeks to uncover the foundational worldview imprinted upon the model. The foundational work of Hofstede on cultural dimensions provides a powerful framework for conceptualizing these differences, such as the spectrum between individualism and collectivism or variations in power distance [hofstede1980cultures]. Recent work has begun to probe these dimensions, asking whether models possess cultural intelligence [bleu2023language] or how they represent knowledge from different cultures [jiang2023large], but a systematic methodology for quantifying these ingrained ”genes” is still nascent. The challenge is immense, as these models are increasingly tasked with sensitive roles in domains like facial expression recognition, where cultural display rules are paramount and label noise can be a significant confounder [chiang2023vicuna].
The stakes of failing to address this challenge are substantial. Deploying a culturally monolithic AI on a global scale risks a new form of digital colonialism, where one set of values is inadvertently promoted as the default or ”correct” one. This can lead to misunderstandings, perpetuate harmful stereotypes [mhamdi-etal-2023-cross], and create systems that are ineffective or even offensive in different cultural contexts. The problem is exacerbated in multimodal systems, where visual and linguistic cues are intertwined, such as in visual dialog or video grounding , and in security applications where understanding human behavior is key, from recognizing micro-actions [ouyang2022training] to detecting malicious intent in side-channel attacks [bai2022training] or against authentication protocols [zhang2022opt].
@@ 2.1 Bias and Fairness in Language Models
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones . Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 2.2 Cross-Cultural and Multilingual NLP
The globalization of AI has spurred a growing interest in cross-cultural and multilingual NLP. Researchers have begun to move beyond English-centric benchmarks [raji2021ai], developing resources for evaluating commonsense reasoning in multiple languages [lin-etal-2021-common]. This has revealed significant disparities, highlighting the poor performance of mainstream models on tasks centered around non-Western cultures, such as those in the Africentric NLP space [ogueji2021state]. The very notion of ”cultural intelligence” in LLMs is now being formally investigated [bleu2023language], with studies directly probing models for cross-cultural differences in values [arora-etal-2023-probing].
@@ 2.3 Probing and Interpreting Model Knowledge
Researchers are now pushing beyond factual knowledge to probe more abstract concepts. Studies have analyzed how models represent social relations and roles [artetxe-etal-2023-representation], and groundbreaking work is using the geometry of a model’s internal activation space to uncover its true ”beliefs,” independent of its textual output [marks2023geometry]. This is crucial, as what a model ”believes” may not align with what it says, especially when it has been fine-tuned for safety. Analyzing the internal mechanisms, such as the structure of attention heads [vig2019analyzing] or using causal analysis to understand syntactic agreement [vig-etal-2020-causal], provides a deeper view into how these representations are formed. The ultimate goal is to understand not just *that* a model holds certain cultural values, but *why* and *how* it applies them, a question that is now being explored in the context of social norms [zhu2024can]. This deep level of analysis is essential for building trustworthy AI, whether it is used for video grounding or for complex, decentralized tasks like federated learning in edge networks .
@@ 2.4 Value Alignment and AI Ethics
This has led to calls for a more critical, decolonizing perspective on digital technologies [mhlambi2023decolonising] and the development of new evaluation frameworks that go beyond standard accuracy metrics. Researchers are now assessing LLMs against international human rights norms [shelby2023norms], testing them for social intelligence [sap-etal-2019-social], and even evaluating their ”personality” using standardized psychological tests [miotto2024do]. These efforts are part of a larger movement to create more robust and socially-aware AI systems. This is critical as AI is deployed in high-stakes, human-centric applications, from developing more efficient model compression techniques and robust neural architecture search methods , to securing our digital lives against sophisticated attacks on authentication systems or privacy via side-channels [hoffmann2022training]. Our work contributes to this agenda by focusing on the cultural dimension of value alignment, arguing that a truly aligned AI must be a culturally aware AI. While benchmarks for Polish politics [kocon2023arrabest] or heterogeneous information retrieval [thakur2021beir] are steps in the right direction, a systematic methodology for comparing foundational cultural values is a critical, and currently missing, piece of the puzzle.
@@ 3.1 Bias and Fairness in Language Models
However, much of this foundational work implicitly focuses on biases within a single, predominantly Anglo-American, cultural context. While critical, it often overlooks the more nuanced, value-laden biases that differ across cultures. Furthermore, research has shown that many debiasing techniques may only offer a superficial fix, covering up systematic biases without truly removing them from the model’s underlying geometric representations [gonen-goldberg-2019-lipstick]. This suggests that a deeper understanding of how biases are formed and propagated is necessary. The trail of political biases, for instance, can be tracked from pretraining data all the way to downstream task performance [feng-etal-2023-pretraining], indicating that the data itself is a primary vector for ideological imprinting. The challenge is amplified in complex, real-world applications where data quality is paramount, such as in developing robust facial expression recognition systems that must contend with noisy labels, a problem that itself can have cultural undertones . Understanding and mitigating these biases is a complex process [liang-etal-2021-towards], requiring a shift from merely detecting bias to understanding its cultural origins.
@@ 5.1.2 Cultural Alignment Index (CAI)
The Cultural Alignment Index (CAI), as defined in Equation, measures how closely each model’s performance aligns with the established Hofstede scores for the USA (a high-IDV, low-PDI country) and China (a low-IDV, high-PDI country). The results are unambiguous. For the IDV dimension, $M_{W}$’s alignment with the USA was 0.91, while its alignment with China was only 0.45. The pattern was reversed for $M_{E}$, which showed an alignment of 0.85 with China and 0.48 with the USA. The same cross-alignment was observed for the PDI dimension.
@@ 5.4.2 Implications for Global AI Deployment and Ethics
Therefore, a one-size-fits-all approach to value alignment is untenable [hendrycks2021aligning]. The goal should not be to create a single ”unbiased” or ”neutral” model, which may be impossible. Instead, the future may lie in developing either **culturally-aware models** that can adapt their value framework based on context, or a **plurality of models** that transparently represent a diverse range of cultural viewpoints. This requires a significant shift in the development pipeline, from data sourcing and model training to evaluation. The need for robust systems that can handle heterogeneity is a common thread in modern computing, from federated learning with non-IID data to wireless systems that must operate in noisy, unpredictable environments .
@@ 5.4.3 Limitations and Future Work
Our study, while providing strong evidence, has several limitations that open avenues for future research. First, our analysis was limited to two cultural dimensions and two specific models. Future work should expand this framework to include more dimensions (e.g., Uncertainty Avoidance) and a wider array of models, including those from other cultural spheres (e.g., India, the Middle East). Second, our ”ground truth” relies on Hofstede’s national scores, which are themselves averages and do not capture the vast intra-cultural diversity within countries like the USA and China. Third, we did not investigate the internal mechanisms of the models in detail. Future work could use techniques from mechanistic interpretability [marks2023geometry, vig-etal-2020-causal] to pinpoint exactly where and how these cultural representations are stored in the network. Finally, our study is diagnostic. The next critical step is prescriptive: developing methods to effectively steer or mitigate these ingrained cultural biases, moving beyond simple debiasing techniques [gonen-goldberg-2019-lipstick] to more fundamental interventions in the training and fine-tuning process. This is a grand challenge, akin to building systems that are robust not just to data noise but to the very fabric of cultural bias woven into human language.